OpenStudy (anonymous):

Solve by elimination 9x - 3y = 24 7x - 3y = 20

OpenStudy (anonymous):

@bahrom7893

OpenStudy (bahrom7893):

what did u get?

OpenStudy (anonymous):

the y are the same already

OpenStudy (anonymous):

do i just go or what

OpenStudy (primeralph):

Add.

OpenStudy (bahrom7893):

oh so just find x

OpenStudy (anonymous):

-1*(9x-3x=24) -9x+3x=-24 7x-3y=20

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

-2x=-4 x=2

OpenStudy (anonymous):

dont give answer plz

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

np

OpenStudy (anonymous):

okay its 2 lol

OpenStudy (anonymous):

its like 2.something

OpenStudy (anonymous):

but i rounded it off

OpenStudy (anonymous):

lol

OpenStudy (primeralph):

It's actually just 2

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

is the answer (2,2)

OpenStudy (anonymous):

@bahrom7893 is it right

OpenStudy (bahrom7893):

2,2 is wrong

OpenStudy (anonymous):

what did i do wrong

OpenStudy (bahrom7893):

9x+7x = 24+20 16x=44 x=44/16=11/4

OpenStudy (primeralph):

Yeah, true.

OpenStudy (bahrom7893):

Huh? 9x - 3y = 24 9*2 - 3*2 = 18 - 6 = 12

OpenStudy (bahrom7893):

yea lol sorry, i was confused in the beg.

OpenStudy (bahrom7893):

But I think I'm right so far.. so x = 11/4 9x - 3y = 24 99/4 - 3y = 24

OpenStudy (primeralph):

Yeah.

OpenStudy (bahrom7893):

33/4 - y = 8 33 - 4y = 32 -4y = 32 - 33 = -1 y = 1/4

OpenStudy (anonymous):

so take th 1/4 and plug it in for y right

OpenStudy (bahrom7893):

well that already is the answer. I got 11/4 for x and plugged it in to find y

OpenStudy (anonymous):

o kk thx that was hard